Delta Corporation’s production and sales data provided below:

                                                            Per Unit                      Percentage

            Selling price                           $9.00                            100%

            Variable cost $6.00 66.67%   

            Contribution margin $3.00 33.33%

           

            Total fixed costs                     $30,000

Compute a break-even point in units and dollars if the following changes occurred:

Selling price decreased by one-tenth.

Variable costs decreased by one-third.

Fixed costs increased by one-third.
Answer #1

* Calculation : Revised Per unit amount Selling Price = variable Cost = (9 (9xvio) = 9-0.9 = $8.10 (6-(68\/3) = 6-2 = 84.00 T